# Simple Yet Effective Selective Imputation for Incomplete Multi-view Clustering

### v1

@@ V Conclusion
In this paper, we propose ISMVC, a simple yet effective informativeness-based selective imputation framework for incomplete multi-view clustering. By quantifying information support for each missing position, ISMVC imputes only well-supported positions and models imputation uncertainty within a variational inference framework. This approach strikes a principled balance between leveraging imputation benefits and mitigating potential adverse effects. Additionally, the core Informativeness-Based Selective Imputation (IBSI) module serves as a lightweight, plug-and-play component across diverse clustering paradigms. Extensive experiments demonstrate its effectiveness and generality across diverse datasets and models.

### v2

@@ V Conclusion
In this paper, we propose SI3, a simple yet effective informativeness-based selective imputation method for incomplete multi-view clustering. By quantifying imputation-relevant informativeness at each missing position, SI3 imputes only well-supported positions and models imputation uncertainty within a variational inference framework. This approach strikes a principled balance between leveraging imputation benefits and mitigating potential adverse effects. Additionally, the Informativeness-Based Selection and Indication module functions as a lightweight and plug-in component, enabling seamless integration with a wide range of clustering paradigms. Extensive experiments demonstrate its effectiveness and generality across diverse datasets and baselines.
Acknowledgments
This research was supported by the National Natural Science Foundation of China (Grant Nos. 62133012, 62303366) and Natural Science Basic Research Program of Shaanxi under Grant No.2023-JC-QN-0648.